OpenStudy (anonymous):

find arc length of x=sqrt(y)-y from y= 1 to 4

OpenStudy (anonymous):

Do you know the arc length formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay so what don't you get?

OpenStudy (anonymous):

sqrt1+[f'x]^2

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

\[\int\limits_{1}^{4}.5y ^{-.5}\] is what i get when i do the formula

OpenStudy (anonymous):

i integrate and get \[y^{.5}/.5\]

OpenStudy (anonymous):

forgot the .5, but they cancel out anyways, so your left with just y^.5

OpenStudy (anonymous):

once i plug in my values though, i dont get the answer in my book

OpenStudy (anonymous):

What about the square root?

OpenStudy (anonymous):

the [f'x]^2 cancels the square root in the arc length formula no?

OpenStudy (anonymous):

Okay, first of all, what is your \(f'\)?

OpenStudy (anonymous):

\[f'(y)=.5y ^{-.5}-1\]

OpenStudy (anonymous):

And when you square that, what do you get?

OpenStudy (anonymous):

not entirely sure, i dont square that since the sqrt in the arc lenth formula should cancel the square i have to do

OpenStudy (anonymous):

so it should in the end be 1+(.5y^-.5)-1

OpenStudy (anonymous):

1's cancel out, and i should have left the integral of .5y^-.5

OpenStudy (anonymous):

But it's under the square root.

OpenStudy (anonymous):

The thing is: \(\sqrt{1+[f']^2}\neq \sqrt{1+f'}\)

OpenStudy (anonymous):

\[\sqrt{1+f'(x)^2} = 1+f'(x)\]

OpenStudy (anonymous):

which is what i did

OpenStudy (anonymous):

Nope, that isn't true either!

OpenStudy (anonymous):

Whenn you square your derivative you should get 1/(4y)-1/y^(.5)+1

OpenStudy (anonymous):

yes i get that. but then i have to take the square root of that, so why bother squaring in the first place? ._.

OpenStudy (anonymous):

You need to add 1 to the square of the derivative before you can take the square root of it.

OpenStudy (anonymous):

\[ \sqrt{1+x^2}\neq 1+x \]

OpenStudy (anonymous):

\[ (1+x)^2 = x^2+2x+ 1\neq 1+x^2 \]

OpenStudy (anonymous):

so then how would i go about taking the square root of something that complex? is there a trick? square root of 1/(4y)-1/y^(.5)+2 yh?

OpenStudy (anonymous):

First just expand it and simplify.

OpenStudy (anonymous):

In most cases it creates a perfect square for you to just be able to take the square root of the whole thing. That's not always the case, however.

OpenStudy (anonymous):

not entirely sure i have to expand something with negative exponents and decimal coiefficients

OpenStudy (anonymous):

my prof knew some type of trick, im not entirely sure what it was though. i dont see how anyone can take the square of what we have right now

OpenStudy (anonymous):

What do you have?

OpenStudy (anonymous):

\[\sqrt{.25y ^{-1}-y ^{-.5}+2}\]

OpenStudy (anonymous):

It doesn't come out to a perfect square. In most cases, it does. Do you have the problem down written correctly? There is just one part missing that keeps this from being a perfect squre.

OpenStudy (anonymous):

\[x=\sqrt{y}-y \] evaluate from y =1 to 4 is the problem exactly how it appears in the book

OpenStudy (anonymous):

ok...let me try something. Give me a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is problem number 5 so i dont see how the it can be so complicated ._.

OpenStudy (anonymous):

There are just some that get all twisted. In most cases, you just have to use some algebraic manipulation. I may have something. Hang on.

OpenStudy (anonymous):

Nope, my middle term comes out to be 2/(2y)^(.5). I need a 1/y^(.5).

OpenStudy (anonymous):

I have tried to factor it about three different ways, and each time my middle term comes out a little bit diffferent than what should be there.

OpenStudy (anonymous):

Alright, im going to go ahead and close this. seems to be not worth the effort. thanks for the help wio and calmat

OpenStudy (anonymous):

Yw. I will shoot you an @ if I come up with something. I am determined to figure it out.

OpenStudy (anonymous):

lol alright

OpenStudy (sirm3d):

\[\large {x=\sqrt y - y\\x'=\frac{1}{2\sqrt y} - 1 = \frac{1-2\sqrt y}{2\sqrt y}\\(x')^2=\frac{1-4\sqrt y+4y}{4y}\\1+(x')^2=\frac{1-4\sqrt y+8y}{4y}\\\sqrt{1+(x')^2}=\frac{1}{2\sqrt y}\sqrt{8y-4\sqrt y+1}=\frac{\sqrt 8}{2\sqrt y}\sqrt{(\sqrt y)^2-(1/2)\sqrt y+(1/8)}\\\quad=\frac{\sqrt 8}{2\sqrt y}\sqrt{\left[(\sqrt y)^2-(1/2)\sqrt y+(1/16)\right]+(1/16)}\\\frac{\sqrt 8}{2\sqrt y}\sqrt{\left[\sqrt y - (1/4)\right]^2+(1/16)} }\]

OpenStudy (sirm3d):

let \[\large{\sqrt y - (1/4) = (1/4)\tan \theta\\\frac{dy}{2\sqrt y}=(1/4)\sec^2\theta d\theta\\\sqrt{\left[\sqrt y - (1/4)\right]^2 + (1/16)}=(1/4)\sec \theta}\]

OpenStudy (sirm3d):

\[\large {\sqrt{1+(x')^2}dx=\sqrt 8 \int \sqrt{(\sqrt y)^2-(1/2)\sqrt y+(1/16)}\frac{dy}{2\sqrt y}\\=\sqrt 8 \int (1/4)\tan \theta (1/4)\sec^2\theta d\theta}\]